OpenStudy (anonymous):

Solve each system using substitution. c(n)=-3n-6 c(n)=n-4 PLEASE HELP :(

OpenStudy (the_fizicx99):

@Lena772 is pro at this

OpenStudy (lena772):

lol

OpenStudy (the_fizicx99):

She's in calc!! :O she knows it ;D ♥

OpenStudy (lena772):

c(n)=-2n-10

OpenStudy (anonymous):

I'm so confused, i worked it every way i can and still don't get the right answer

OpenStudy (lena772):

Just add it together.

OpenStudy (lena772):

Ok nvm I'm doing the wrong method

OpenStudy (anonymous):

My choices are,A. -1/4, -3 1/8 B. -3/4, -2 1/2 C. -1/2, -4 1/2

OpenStudy (lena772):

c(n) = -3(n-4) - 6 c(n) = -3n + 12 - 6 c(n) = -3n + 6